Determine, with proof, the supremum and infimum of the set A={(n+1)/n | n ∈ N}
1
Expert's answer
2020-12-22T20:17:08-0500

Consider the set "A=\{\dfrac{n+1}{n}:n\in\N\}"

Since "\dfrac{n+1}{n}=1+\dfrac{1}{n}>1" for each "n\in\N, 1" is a lower bound of the set. Moreover, for each "\varepsilon>0," by Archimedean Property, there exists "N\in\N" such that "N\geq\dfrac{1}{\varepsilon}." Then "\dfrac{1}{N}\leq\varepsilon" and therefore "1+\varepsilon" is not a lower bound of the set for each "\varepsilon>0." We conclude that "1" is the infinum of the set "\{\dfrac{n+1}{n}:n\in\N\}."


Since "n\geq1" we have "\dfrac{n+1}{n}=1+\dfrac{1}{n}\leq1+\dfrac{1}{1}=2, n\in\N"

The set "A" is bounded above: there is "M=2" such that "\dfrac{n+1}{n}\leq2" for all "n\in\N."

The number "M=2" is the upper bound of the set "A."

Since "2\in A" then we conclude that "2" is the supremum of the set "\{\dfrac{n+1}{n}:n\in\N\}."